r Enter all amounts as positive values 10,000 32,100 26,001
Answer #1
Case 1
Manufacturing overhead applied 15000 =10000*150%
Total current manufacturing costs 36000 =11000+10000+15000
Cost of goods manufactured 40000 =8600+36000-4600
Cost of goods sold 34800 =3700+40000-8900
Case 2
Direct materials used 3333 =25000-13000-8667
Direct labor 8667 =13000/150%
Beginning work in Process inventory 32900 =48000+9900-25000
Ending finished goods inventory 20000 =48000+12000-40000
Case 3
Direct labor 9000 =(32100-9600)/250%
Manufacturing overhead applied 13500 =9000*150%
Ending work in Process inventory 14599 =8500+32100-26001
Beginning finished goods inventory 19799 =39000+6800-26001